OpenStudy (jazzyfa30):

Answer maximum value of ____ or minimum value of (blank) Determine whether each function has a maximum or a minimum value. Then find the maximum or minimum value of each function. f(x) = x2 +9

OpenStudy (jazzyfa30):

@doulikepiecauseidont

OpenStudy (jazzyfa30):

@calculusxy

OpenStudy (calculusxy):

Hi

OpenStudy (jazzyfa30):

hi

OpenStudy (anonymous):

it is a parabola that opens up, so it has a minimum value, no maximum

OpenStudy (anonymous):

since the smallest \(x^2\) can be is zero, the smallest \(x^2+9\) can be is \(9\)

OpenStudy (calculusxy):

Well the 2 is squared since it is after the x.

OpenStudy (jazzyfa30):

ok

OpenStudy (calculusxy):

So as the equation goes the answer will be found in the equation of x^2+9

OpenStudy (calculusxy):

Remember this formula. (X+a)(x+b)= x^2+bx+ax+ab= x^2+(ab)^2+ab

OpenStudy (calculusxy):

Sorry if you get confused.

OpenStudy (jazzyfa30):

confussed i dont understand can you walk me through it please

OpenStudy (anonymous):